OpenStudy (anonymous):

domain of y=x/(x^2-9)

OpenStudy (anonymous):

the domain is a set of numbers. When you get functions like this also pay attention to the denominator. Answer this question: what value for x will give you '9'

OpenStudy (anonymous):

\[y=\frac{ x }{ x^2-9 }\]

OpenStudy (anonymous):

9

OpenStudy (anonymous):

...look at the denominator. For what value of x (if you substitute in a number) will give you 9 where it says x^2? Why am i asking you this? Because the function is U N D E F I N E D if the DENOMINATOR = 0. So with problems like this always solve for the denomintor = 0 and exclude that number(s) from the domain that makes the denominator = to 0. And in this case it will be THREE. Because 3^2 = 9. And 9-9 = 0 !!!!!!!!!!!!! THUS Your domain is \[x \epsilon (-infinity,3)U(3,+infinity)\]

OpenStudy (anonymous):

I gave you the domain in INTERVAL NOTATION. Do you understand interval notation as I have written?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cant u also right x cannot equal 3 or -3

OpenStudy (anonymous):

You were wrong when I asked you your question of what value of x will be 9 in the DENOMINATOR (Remember, we don't care about the numerator. Because it will be valid for any number of x.) Simply rewite the denominator in cases like this and solve for 0. That x value that makes the denominator 0 (in this case 3) is the value that x CANNOT BE. WHY!??? BECAUSE it will make the function U N D E F I N E D -------> If you divid any number by 0 you get an ERROR / UNDEFINED. you understand?

OpenStudy (anonymous):

Yes, you can also write it like x=/= 3 or x=/=-3.

OpenStudy (anonymous):

I know I was wrong im asking can I also write x cannot equal -3 or 3 instead of wat u wrote? yes or no question

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

now I took the first derivative and need to find critical numbers

OpenStudy (anonymous):

i set x^2-9=0 and i got -3 and +3 but since there not in the domain i don't have critical numbers right?

OpenStudy (anonymous):

have you solved for the FIRST DERIVATIVE of the whole function yet?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then i set x^2-9=0 and i got -3 and +3 but since there not in the domain i don't have critical numbers right?

OpenStudy (anonymous):

thats not correct.

OpenStudy (anonymous):

i got 1/x^2-9 for first derivative

OpenStudy (anonymous):

post dy/dx (the derivative function) for me please.

OpenStudy (anonymous):

Oh, is THAT the first derivative?

OpenStudy (anonymous):

yes that's the answer to my first derivative

OpenStudy (anonymous):

\[dy/dx(x/x^2-9) = 1/x2-9\]

OpenStudy (anonymous):

ok, what you need to do is the following: 1.) Solve for f'(x) = 0 or dy/dx = 0) Do you know how to do that?

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

no, you solved for the denominator = 0. Thats for the DOMAIN.

OpenStudy (anonymous):

c=+-3

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so it would just be 1=0

OpenStudy (anonymous):

You need to solve for the DERIVATIVE FUNCTION = 0. CRITICAL POINTS are those points in which the GRADIENT (Derivative) of a function is =0.

OpenStudy (anonymous):

0=1/x^2-9

OpenStudy (anonymous):

so there are no critical numbers right

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

i have another question too

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

y=x/(x^2-9) for y to be defined x^2-9\[\neq 0\],x^2\[\neq9\] x≠3 x<-3 or x>3 combining domain is (-∞,-3)∪(3,∞)

OpenStudy (anonymous):

thank u :)

OpenStudy (anonymous):

the critical points are -3 and 3, is that what you got?